OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 2x + 5y = 4? (3, –2) (7, –2) (0, –1) (0, 1)

OpenStudy (solomonzelman):

In other words, which of the pairs is a solution to the equation. 2x + 5y = 4 2(3) + 5(-2) = 4 ---> 6-10=4 Not A. 2(7) + 5(-2) = 4 ---> 14-10=4 YES B. 2(0) + 5(-1) = 4 ---> -5=4 Not C. 2(0) + 5(1) = 4 ---> 5=4 Not D.

OpenStudy (solomonzelman):

I plugged in the values, that's all.

OpenStudy (anonymous):

ok i have a question there is one im really confused on may u help

OpenStudy (solomonzelman):

Sure.

OpenStudy (anonymous):

theres D

OpenStudy (solomonzelman):

\[y-4=2(x+1)\]set it into a slope intercept formula, or into a y=mx+b form. \[y-4=2x+2\]\[y=2x+6\] choose any random value for x. lets say x=-2 y=2(-2)+6 ---> y=2, so your first point is (-2,2) Now, lets say x=-3 y=2(-3)+6 ---> y=0, so your first point is (-3,0) also the always given y-intercept b, which is (0,6) which line has points (-2,2) , (-3,0) and (0,6) ?

OpenStudy (solomonzelman):

2nd graph on the second attachment.

OpenStudy (anonymous):

well my slope says 0,6 which im getting from the intercept of the last one D

OpenStudy (solomonzelman):

your slope says? I know that slope is 2, and y-intercept is 6 agree?

OpenStudy (anonymous):

yes i completely agree but looking back on all of my ther ones i always tend to mix the 2 up

OpenStudy (anonymous):

What up @SolomonZelman